# IOI: Invisible One-Iteration Adversarial Attack on No-Reference Image- and Video-Quality Metrics

### v1

@@ 6 Discussion
Different frame frequency and attack success. We conducted additional experiments to show the importance of a one-iteration setup when attacking NR quality metrics in videos. In Appendix B, we show that one iteration for each frame is more effective than $k$ iterations for each $k$-th frame. Table 4 summarizes the results. The reason lies in the difference of quality metrics application to videos from the classification task, where an attacker can affect only several frames to fool the classifier.
k=1 k=2 k=4 k=6 k=8 k=10
RG 13.2% 10.0% 6.4% 4.8% 3.9% 3.1%
Table 4: Results of attacking PaQ-2-PiQ (Ying et al., 2020) on one video on each $k$-th frame for k iteration.
IOI performance under defences. We made additional experiments (Table 5) to check the robustness of the proposed method to three adversarial defences: video compression (Shaham et al., 2018), random crop and resize used in (Shumitskaya et al., 2023a) for NR metrics. Defences were evaluated for videos from the DERF dataset. Although the proposed method affects only high-frequency information, video compression reduced relative gain only by 2.4%. Random cropping confuses the attack and reduces relative gain approximately two times. Frames resizing almost completely mitigates the relative gain from 14.6% to 1%; however, an NR metric increase by 1% is still significant for benchmarks; sometimes, teams compete to achieve a 0.1% metric increase to win the competition. More details are presented in Appendix D.
@@
Table 5: IOI performance under adversarial purification defences (video compression, random cropping and resizing). Relative gain averaged for 12 videos.
Speed of the proposed method. The PyTorch realization of the IOI attack allows reaching 8 fps on the NVIDIA Tesla T4 GPU. Details presented in Appendix B.1.
Limitations. Our method works in a white-box scenario that implies that an attacker knows and has access to the target model. The white-box scenario is less universal than a black-box; however, as described in the introduction, NR quality metrics are usually published as part of the benchmark methodology. We do not compare our attack with black-box attacks because of the significant difference in operation speed; black-box attacks are unlikely to be injected into video processing algorithms that compete in quality and speed, which is the scenario we target in our work. We considered only NR metrics in this research, as FR metrics are much more difficult to attack in real-life scenarios. The robustness of FR metrics has been studied in (Ghildyal & Liu, 2023).
@@ 7 Conclusion
This paper introduces the IOI adversarial attack on NR image- and video-quality metrics. Its primary objective is to generate imperceptible perturbations for images or videos using only one iteration. Through extensive experiments, we showed that existing methods fail to produce high-fidelity adversarial videos in near real-time scenarios (1 – 10 fps). In contrast, our proposed method demonstrates better effectiveness at high speed. Subjective and objective comparisons showed that the proposed method produces adversarial images and videos of superior visual quality, achieving the same attack success and speed as prior methods. The proposed attack is a potent tool for experimentally assessing the vulnerability of NR quality metrics. By publishing our method, we provide a tool for verification of NR metrics robustness for benchmark organizers and contribute to the future development of robust image- and video-quality metrics. The proposed method can be used as a part of an adversarial training technique to improve the robustness of image- and video-quality metrics. Our code is openly accessible at link is hidden for blind review.
Impact Statements
@@ Appendix A Proof of Theorem 1
Theorem 1. Let $I\text{ and }I^{p}$ be original and perturbed image correspondingly, $I^{a}$ – IOI adversarial image based on $I^{p}$ with truncating parameter $f$. Then inequality 12 is correct, where $MAE^{*}(\cdot,\cdot)$ is given by Equation 3.
@@
Appendix B Attack speed and efficiency at different frame frequencies
This section demonstrates that employing a single iteration for each frame produces superior results compared to the sporadic application of multiple iterations, such as ten iterations for every tenth frame. To validate this hypothesis, we conducted an additional experiment. We selected the PaQ-2-PiQ (2020) NR metric, “Controlled Burn” video from the DERF dataset (der, 2001), and applied the I-FGSM attack (Kurakin et al., 2018). I-FGSM is an extension of FGSM involving multiple iterations and is the foundation for more advanced adversarial attacks.
We executed I-FGSM for different iteration counts: $n=1$, $n=2$, $n=4$, $n=6$, $n=8$, and $n=10$. In each experiment, only a fraction of frames underwent attack, specifically $\frac{1}{n}$. The selection of frames for the attack was done uniformly. The results of these experiments are illustrated in Figure 4. As the number of iterations in the attack increases, there is a corresponding rise in the attack’s relative gain on each particular frame. However, evaluating the overall relative gain involves averaging relative gains across all frames. The optimal averaged relative gain occurs when each frame is attacked with just one iteration, and this gain decreases monotonically with the increase in the parameter $n$. We also measured the computation time for attacks with different values of $n$, which remained nearly constant at 3 seconds for all attacks. This consistency arises from the fact that each experiment’s total number of adversarial iterations was the same.
In conclusion, attacking each frame with just one iteration proves more effective than employing $k$ iterations to attack every $k$-th frame. This conclusion is drawn from the observation that, compared with other values of $n$, a one-iteration attack yields superior averaged relative gain within the same attack time. This highlights that attacks on video quality metrics differ from the classification task, where an attacker can affect only several frames to fool the classifier, and emphasizes the importance of one-iteration attack approaches. From the results of this experiment, we can conclude that the effectiveness of adversarial attacks for video quality metrics is defined by its effectiveness at one-iteration setup.
B.1 Speed for one iteration
@@
IOI (ours) ${}_{PyTorch\text{ }realization}$ 0.028 sec 7.81 fps
@@ Appendix C Experiment with multiple iterations
We conducted an additional experiment to evaluate the performance of visual-oriented methods employed in this paper in the context of multiple iterations.
For $n=10$ and $n=20$ iterations, we run the proposed method with $\epsilon=0.1$ and step size of $\frac{2\epsilon}{n}$. These experiments used the PaQ-2-PiQ NR model (Ying et al., 2020); the relative gain was 13% for both 10 and 20 iterations. Utilizing Algorithm 1, we searched for the minimal $lr$ parameter in other attacks to achieve the same relative gain. Subsequently, we evaluated the visual quality of the resulting adversarial images using four FR metrics: SSIM, PSNR, VIF, and LPIPS. The results for $n=10$ and $n=20$ iterations are presented in Tables 7 and 8 respectively. At 20 iterations, all methods produce nearly identical results, highlighting the primary strength of the proposed IOI method in its effectiveness in one-iteration settings.
@@
Appendix D IOI performance under defences
We conducted additional experiments to evaluate the robustness of the proposed IOI attack to three defences: compression (Shaham et al., 2018), random crop and resize (Shumitskaya et al., 2023a).
@@ Appendix D IOI performance under defences
Random crop and resize defences were previously studied in (Shumitskaya et al., 2023a) for evaluation of UAP (Shumitskaya et al., 2022) against NR quality metrics. Authors showed that random crop and resize to 80% of the original image/video size help to improve NR quality metric robustness to adversarial attacks without significant loss in correlations with subjective scores. Because of that, we chose parameter 80% for our experiments. Defended transformation in the case of random crop defence was selecting a random crop of the frame with 80% of the original frame size. An image was resized to 80% of the original frame size for resize-based defence. Shaham et al. showed that compression can be used as defence (Shaham et al., 2018). In our experiment, defended transformation for compression defence was compression with a CRF value 16 using x264 video codec (preset “Medium”).
@@
For the experiment, we used 12 adversarial videos generated against PaQ-2-PiQ (source videos from the DERF dataset (der, 2001)) and original ones. We measured relative gain for the IOI attack with and without defences. For relative gain measurement under defence, we calculated PaQ-2-PiQ scores for original and adversarial videos after defended transformation and then, based on these scores, calculated relative gain. Results are presented in the Table 9. Compression defence reduced relative gain only by 2.4%. Random crop confuse attack and reduce relative gain approximately two times. Resize defence almost completely mitigate the relative gain, however gain in 1% still can be significant in benchmarks.
From the results of these experiments, we can conclude that the proposed IOI attack is robust to compression and random crop defences but vulnerable to the resize defence.
@@
Appendix E Video sequences

### v2

@@ 6 Discussion
Different frame frequency and attack success. We conducted additional experiments to show the importance of a one-iteration setup when attacking NR quality metrics for videos. This section demonstrates that employing a single iteration for each frame produces superior results compared to the sporadic application of multiple iterations, such as ten iterations for every tenth frame. We selected the PaQ-2-PiQ (2020) NR metric, “Controlled Burn” video from the DERF dataset (2001), and applied the I-FGSM attack (Kurakin et al., 2018). I-FGSM is an extension of FGSM, involving multiple iterations.
We executed I-FGSM for different iteration counts: $n=1$, $n=2$, $n=4$, $n=6$, $n=8$, and $n=10$. Only some frames underwent attack in each experiment, specifically $\frac{1}{n}$. The selection of frames for the attack was done uniformly. The results of these experiments are illustrated in Figure 4. As the number of iterations in the attack increases, there is a corresponding rise in the attack’s relative gain on each particular frame. However, evaluating the overall relative gain involves averaging relative gains across all frames. The optimal averaged relative gain occurs when each frame is attacked with just one iteration, and this gain decreases monotonically with the increase in the parameter $n$. We also measured the computation time for attacks with different values of $n$, which remained nearly constant at 3 seconds for all attacks. This consistency arises from the fact that each experiment’s total number of adversarial iterations was the same.
Compared with other values of $n$, we can see that a one-iteration attack yields superior averaged relative gain within the same attack time. This highlights that attack on video quality metrics differ from the classification task, where an attacker can affect only several frames to fool the classifier. From the results of this experiment, we can conclude that the effectiveness of adversarial attacks for video quality metrics is defined by their effectiveness at a one-iteration setup.
Speed of the proposed method. The PyTorch realization of the IOI attack allows reaching 8 fps on the NVIDIA Tesla T4 GPU. Details presented in Appendix B.
IOI performance under defences. We did additional experiments (Table 4) to check the robustness of the proposed method to three adversarial defences: video compression (Shaham et al., 2018), random crop and resize used in (Shumitskaya et al., 2023a) for NR metrics. Defences were evaluated for videos from the DERF dataset. Although the proposed method affects only high-frequency information, video compression reduced relative gain only by 2.4%. Random cropping confuses the attack and reduces relative gain approximately two times. Frames resizing almost completely mitigates the relative gain from 14.6% to 1%; however, an NR metric increase by 1% is still significant for benchmarks; sometimes, teams compete to achieve a 0.1% metric increase to win the competition. More details are presented in the Appendix E.
@@
Table 4: IOI performance under adversarial purification defences (video compression, random cropping, and resizing). Relative gain averaged for 12 videos.
Limitations. Our method works in a white-box scenario that implies that an attacker knows and has access to the target model. The white-box scenario is less universal than a black-box; however, as described in the introduction, NR quality metrics are usually published as part of the benchmark methodology. We made additional experiments of analysing black-box transferability (Appendix F) and IOI performance in black-box settings (Appendix G). We found out low transferability across different models and significant difference in operation speed of white-box and black-box attacks. Black-box attacks are unlikely to be injected into video processing algorithms that compete in quality and speed, which is the scenario we target in our work. We considered only NR metrics, as FR metrics are much more difficult to attack in real-life scenarios. The robustness of FR metrics has been studied in (Ghildyal & Liu, 2023).
Additional experiments. We made the following additional experiments: experiment to compare the proposed method with prior methods when applying different parameters (Appendix D), metric score decreasing experiment (Appendix H), IOI attack on segment-level video quality model (Appendix I). We also measured time spending for one-iteration for all methods tested in this paper (Appendix B).
@@ 7 Conclusion
This paper introduces the IOI adversarial attack on NR image- and video-quality metrics. Its primary objective is to generate imperceptible perturbations for images or videos using only one iteration. Through extensive experiments, we showed that existing methods fail to produce high-fidelity adversarial videos in near real-time scenarios (1 – 10 fps). In contrast, our proposed method demonstrates better effectiveness at high speed. Subjective and objective comparisons showed that the proposed method produces adversarial images and videos of superior visual quality, achieving the same attack success and speed as prior methods. The proposed attack is a potent tool for experimentally assessing the vulnerability of NR quality metrics. By publishing our method, we provide a tool for verification of NR metrics robustness for benchmark organizers and contribute to the future development of robust image- and video-quality metrics. The proposed method can be used as a part of an adversarial training technique to improve the robustness of image- and video-quality metrics. Our code is openly accessible at https://github.com/katiashh/ioi-attack.
Acknowledgements
The authors would like to thank the CMC Faculty of MSU, especially A.V. Gulyaev for providing the necessary computing resources, which enabled us to undertake some of the calculations for this paper. We also would like to express our gratitude to Mikhail Pautov for discussing the results of this research and assistance in theoretical statements.
The work was supported by a grant for research centers in the field of artificial intelligence, provided by the Analytical Center in accordance with the subsidy agreement (agreement identifier  [^00] D730321P5Q0002) and the agreement with the Ivannikov Institute for System Programming of dated November 2, 2021 No. 70-2021-00142.
Impact Statement
@@ Appendix A Proof of Theorem 1
Theorem 1. Let $I\text{ and }I^{p}$ be original and perturbed image correspondingly, $I^{a}$ – IOI adversarial image based on $I^{p}$ with truncating parameter $f$. Then inequality 13 is correct, where $MAE^{*}(\cdot,\cdot)$ is given by Equation 3.
@@
Appendix B Speed for one iteration
@@
IOI (ours) PyTorch 0.028 sec 7.81 fps
@@ Appendix C Experiment with multiple iterations
We conducted an additional experiment to evaluate the performance of visual-oriented methods employed in this paper in the context of multiple iterations. For $n=10$ and $n=20$ iterations, we run the proposed method with $\epsilon=0.1$ and step size of $\frac{2\epsilon}{n}$. These experiments used the PaQ-2-PiQ NR model (Ying et al., 2020); the relative gain was 13% for both 10 and 20 iterations. Utilizing Algorithm 1, we searched for the minimal $lr$ parameter in other attacks to achieve the same relative gain. Subsequently, we evaluated the visual quality of the resulting adversarial images using four FR metrics: SSIM, PSNR, VIF, and LPIPS. The results for $n=10$ and $n=20$ iterations are presented in Tables 6 and 7 respectively. At 20 iterations, all methods produce nearly identical results, highlighting the primary strength of the proposed IOI method in its effectiveness in one-iteration settings.
@@
Appendix D Different parameter’s comparison
We made an additional experiment to compare the proposed method with prior methods when applying different parameters. This allows us to compare methods in slightly different attack strengths. Results in the Table 8 showed that the proposed method generates images with better visual quality when achieving the same increase in metric score for three different increase levels.
Table 8: Experiment of methods comparison under different relative gains, corresponding to different $\epsilon$ levels in the proposed IOI method.
$\epsilon$ Method SSIM $\uparrow$ PSNR $\uparrow$ VIF $\uparrow$ LPIPS $\downarrow$
0.08 FGSM (2015), SSAH (2022), Zhang et al. (2022b) 0.934$\pm$0.006 36.4$\pm$0.2 0.733$\pm$0.009 0.082$\pm$0.008
NVW (2021) 0.940$\pm$0.006 37.4$\pm$0.4 0.745$\pm$0.009 0.072$\pm$0.007
Korhonen et al. (2022b) 0.932$\pm$0.005 36.2$\pm$0.2 0.727$\pm$0.009 0.083$\pm$0.007
AdvJND (2020) 0.812$\pm$0.005 31.9$\pm$0.1 0.466$\pm$0.007 0.151$\pm$0.006
UAP (2022) 0.792$\pm$0.004 27.9$\pm$0.2 0.432$\pm$0.005 0.251$\pm$0.004
FACPA (2023b) 0.888$\pm$0.003 31.7$\pm$0.2 0.586$\pm$0.005 0.151$\pm$0.003
IOI (ours) 0.966$\pm$0.002 35.5$\pm$0.1 0.786$\pm$0.004 0.037$\pm$0.002
0.1 FGSM (2015), SSAH (2022), Zhang et al. (2022b) 0.884$\pm$0.007 33.6$\pm$0.3 0.635$\pm$0.010 0.134$\pm$0.009
NVW (2021) 0.897$\pm$0.007 34.7$\pm$0.5 0.648$\pm$0.011 0.120$\pm$0.008
AdvJND (2020) 0.740$\pm$0.008 29.5$\pm$0.2 0.384$\pm$0.008 0.208$\pm$0.007
UAP (2022) 0.737$\pm$0.004 26.3$\pm$0.2 0.371$\pm$0.004 0.314$\pm$0.005
FACPA (2023b) 0.863$\pm$0.003 30.5$\pm$0.2 0.539$\pm$0.005 0.182$\pm$0.004
IOI (ours) 0.950$\pm$0.002 33.4$\pm$0.2 0.695$\pm$0.005 0.059$\pm$0.003
0.12 FGSM (2015), SSAH (2022), Zhang et al. (2022b) 0.789$\pm$0.016 30.7$\pm$0.5 0.548$\pm$0.015 0.274$\pm$0.022
NVW (2021) 0.795$\pm$0.016 31.5$\pm$0.6 0.555$\pm$0.016 0.264$\pm$0.022
Korhonen et al. (2022b) 0.774$\pm$0.016 30.0$\pm$0.5 0.524$\pm$0.016 0.295$\pm$0.022
AdvJND (2020) 0.696$\pm$0.007 28.4$\pm$0.1 0.340$\pm$0.006 0.239$\pm$0.008
UAP (2022) 0.705$\pm$0.004 25.4$\pm$0.1 0.342$\pm$0.004 0.348$\pm$0.007
FACPA (2023b) 0.758$\pm$0.003 26.9$\pm$0.2 0.392$\pm$0.004 0.291$\pm$0.005
IOI (ours) 0.936$\pm$0.003 32.2$\pm$0.2 0.681$\pm$0.005 0.077$\pm$0.004
Appendix E IOI performance under defences
@@
We conducted additional experiments to evaluate the robustness of the proposed IOI attack to three defences: compression (Shaham et al., 2018), random crop and resize (Shumitskaya et al., 2023a). Random crop and resize defences were previously studied in (Shumitskaya et al., 2023a) for evaluation of UAP (Shumitskaya et al., 2022) against NR quality metrics. Authors showed that random crop and resize to 80% of the original image/video size help to improve NR quality metric robustness to adversarial attacks without significant loss in correlations with subjective scores. Because of that, we chose parameter 80% for our experiments. Defended transformation in the case of random crop defence was selecting a random crop of the frame with 80% of the original frame size. An image was resized to 80% of the original frame size for resize-based defence. Shaham et al. showed that compression can be used as defence (Shaham et al., 2018). In our experiment, defended transformation for compression defence was compression with a CRF value 16 using x264 video codec (preset “Medium”).
@@
For the experiment, we used 12 adversarial videos generated against PaQ-2-PiQ (source videos from the DERF dataset (der, 2001)) and original ones. We measured relative gain for the IOI attack with and without defences. For relative gain measurement under defence, we calculated PaQ-2-PiQ scores for original and adversarial videos after defended transformation and then, based on these scores, calculated relative gain. Results are presented in the Table 9. Compression defence reduced relative gain only by 2.4%. Random crop confuse attack and reduce relative gain approximately two times. Resize defence almost completely mitigate the relative gain, however gain in 1% still can be significant in benchmarks. From the results of these experiments, we can conclude that the proposed IOI attack is robust to compression and random crop defences but vulnerable to the resize defence.
@@
Appendix F Black-box transferability
We conducted an additional experiment to analyse the applicability of IOI in transferable black-box setting. The experiment was organised as follows: for all generated adversarial images from NIPS2017 dataset, we measured PaQ-2-PiQ, Hyper-IQA and TReS, i.e. for adversarial images created to attack PaQ-2-PiQ we also measured quality scores by Hyper-IQA and TReS. Based on these metrics scores, we calculated relative gains. Results are presented in the Table 10. All eight methods tested in the paper showed low transferability to unseen models. Low transferability can be explained by these metrics having completely different architectures: PaQ-2-PiQ employs RoIPool layers, HyperIQA utilizes ResNet50 and TReS is based on transformer architecture. Also, it’s important to note that improving transferability for image quality models is more challenging than for classifiers or detectors. Transferability can occur in classification and detection tasks because different classifiers ”look” at the same regions of images where classified objects are located (Meng et al., 2019). In contrast, different image quality metrics can look at different regions to estimate the score, which inherently complicates the achievement of transferability. Thus, we can infer that developing imperceptible and, at the same time, transferable one-iteration attacks on video-quality models is a challenging problem that we will consider for further research.
Table 10: Results of experiment on transferability of methods used in the paper in one-iteration setting.
Attack PaQ-2-PiQ Hyper-IQA TReS
Test PaQ-2-PiQ Hyper-IQA TReS PaQ-2-PiQ Hyper-IQA TReS PaQ-2-PiQ Hyper-IQA TReS
FGSM, SSAH, Zhang 7.40% -8.76% -14.93% 0.16% 1.03% -18.99% 0.01% -0.39% 4.07%
NVW 7.17% -8.27% -14.19% 0.13% 0.70% -15.32% 0.08% 0.09% 4.50%
Korhonen 7.36% -8.85% -14.81% 0.23% 0.42% -15.97% 0.07% 0.29% 5.04%
AdvJND 6.61% -0.57% -19.84% -0.18% 4.09% -3.96% 0.04% -0.43% 7.02%
UAP 11.62% -6.03% -15.55% 3.13% 4.83% 6.65% 2.44% -0.52% 0.85%
FACPA 9.68% -5.01% -9.14% -0.30% 1.48% 9.52% -6.02% -0.51% 1.14%
IOI (ours) 7.33% -3.56% -6.25% 0.26% 7.86% -1.73% 0.39% 2.34% 7.40%
Appendix G IOI performance in black-box setting
We conducted an additional experiment aimed to show that the proposed method can be adapted for use in a black-box setting. To do this, we replaced an FGSM-generated perturbation with a black-box-generated perturbation, e.g. Square Attack (Andriushchenko et al., 2020). We will call this modification BB-IOI. It’s important to note that such modification transforms the method into a multi-iteration. To verify the efficiency of BB-IOI, we conducted an additional experiment involving the implementation of the BB-IOI attack that consists of two stages: 1) generating the adversarial perturbation in a black-box manner using the Square Attack (Andriushchenko et al., 2020) method and 2) processing this perturbation using frequency and weighting modules. As a result, BB-IOI provides high-quality adversarial images by objective metrics. Although the property of imperceptibility remains, it’s important to note that the computation complexity of the method has increased, which is common for black-box methods. For attacking PaQ-2-PiQ on the NIPS2017 dataset, BB-IOI achieved an average 1.21% gain operating at 40 seconds per image. Results are presented in the Table 11. This performance is 6.12% lower and 1400 times slower than IOI. Furthermore, the proposed IOI method can be adapted for use in transfer-based settings by combining it with transfer-based perturbation generation techniques (Long et al., 2022), (Li et al., 2023).
Table 11: Comparison of the proposed method used in white-box setting (IOI) and black-box setting (BB-IOI).
Method SSIM $\uparrow$ PSNR $\uparrow$ VIF $\uparrow$ LPIPS $\downarrow$ Relative gain $\uparrow$ Time on one image $\downarrow$
IOI 0.945$\pm$0.002 33.4$\pm$0.2 0.756$\pm$0.005 0.059$\pm$0.003 7.33% 0.028 sec
BB-IOI 0.988$\pm$0.001 36.9$\pm$0.1 0.815$\pm$0.003 0.024$\pm$0.001 1.21% 40 sec
Appendix H Metric score decreasing experiment
In this section, we show the possibility of metrics score decreasing. Given that the proposed method consists of two parts (generating a perturbed image using FGSM and subsequent processing in frequency and weighting modules), changing the optimization direction in FGSM leads to guiding an attack in the opposite direction. We conducted additional experiments to show that the proposed method can decrease PaQ-2-PiQ metric scores. We applied the proposed method to attack it on the same NIPS2017 dataset we used to increase this metric. The results revealed that increasing metric scores yielded a 7.32% score increase, while decreasing metric scores resulted in a 7.61% score decrease (almost the same). However, it’s worth noting that we focused on increasing metrics’ scores because decreasing quality metrics’ scores holds less practical significance. An attacker can decrease the metrics for quality camouflage (2020), and it is the only real-life scenario known to the authors.
Appendix I IOI attack on complex VQA metric
The proposed method applies to segment-level video quality models. We conducted an additional experiment targeting the segment-level VQA metric MDTVSFA (Li et al., 2021) to show this. To apply the proposed attack, we first slightly modified MDTVSFA to get access to its gradient (we removed $torch.no\_grad()$ context-manager from the feature extraction module and modified the forward process in the inference model to process batches rather than dictionaries). Then, we applied IOI to attack three videos on a per-frame basis, processing one frame at a time. Subsequently, we constructed adversarial videos from these frames and calculated quality scores using the original segment-level MDTVSFA on these videos. Results are presented in the Table 12. Remarkably, this approach yielded a significant relative gain, with a 15% increase in scores. This experiment showed that attacking only spatial features of a VQA metric without accounting for temporal and other features is enough to achieve high attack success.
Table 12: Experiments of IOI attack targeting segment-level video-quality metric MDTVSFA. The attack was performed per-frame. Resulting gain was calculated using original segment-level MDTVSFA.
Video MDTVSFA clean MDTVSFA IOI attacked
Blue Sky 0.544 0.659 ($\uparrow$11.5%)
Crowd Run 0.555 0.759 ($\uparrow$20.4%)
Pedestrian Area 0.584 0.737 ($\uparrow$15.3%)
Appendix J Video sequences